There were no solutions available for the Part 4 Integrative Case from the "Fundamentals of Corporate Finance" 4th Edition by Berk, DeMarzo and Harford. Could anyone help with solving the Case Questions below based on the information provided?

Base Information:

1. The risk-free rate of interest, in this case, the yield of the 10-year government bond, which is 3%.

2. HydroTech's:

a. Market capitalization (its market value of equity), $100 million.

b. CAPM beta, 1.2

c. Total book value of debt outstanding, $50 million.

d. Cash, $10 million.

3. The cost of debt (using the quoted yields on Hydro Tech's outstanding bond issues), which is 5%.

With this information in hand, you are now prepared to undertake the analysis.

Case Questions:

1. Calculate HydroTech's net debt.

2. Compute HydroTech's equity and (net) debt weights based on the market value of equity and the book value of net debt.

3. Calculate the cost of equity capital using CAPM, assuming a market risk premium of 5%.

4. Using a tax rate of 35%, calculate HydroTech's effective cost of debt capital.

5. Calculate HydroTech's WACC.

6. When is it appropriate to use this WACC to evaluate a new project?
Answer #1

(1) Cash = $ 10 million, Book Value of Debt = $ 50 million

Net Debt = Book Value of Debt - Cash = 50 - 10 = $ 40 million

(2) Market Value of Equity = $ 100 million and Net Debt = $ 40 million

Equity Weight = [100 / (100+40)] = 5/7 and Net Debt Weight = [40 / 140] = 2/7

(3) Market Risk Premium = RPm = 5 %, Beta = 1.2 and Risk-Free Rate = Rf = 3 %

Using CAPM, Equity Cost of Capital = 3 + 1.2 x 5 = 9 %

(4) Tax Rate = 35 % and Quoted Yield on Outstanding Bond Issue = 5 %

Effective Cost of Debt Capital = (1-0.35) x 5 = 3.25 %

(5) WACC = Weight of Debt x Effective Cost of Debt Capital + Weight of Equity x Equity Cost of Capital = (2/7) x 3.25 + (5/7) x 9 = 7.35714 % ~ 7.36 %

(6) It is appropriate to use the firm's WACC for evaluating new projects only when the project(s) under evaluation has a risk-level equal to the overall risk level of the firm.